OpenStudy (anonymous):

What is x in (2x+10)(2x+6)= 96 meters squared

OpenStudy (anonymous):

It's factoring and I usually know how to factor but I'm stumped

OpenStudy (anonymous):

Distribute it properly: \[(a+b)(c+d) = ac + ad + bc + bd\]

OpenStudy (anonymous):

I did that. and then I got 4x squared + 32x +60 and then that equals to 96 meters squared, but then I get stumped after

OpenStudy (anonymous):

2x+6=t 2x+10=t+4 t(t+4)=96 t^2+4t=96 t^2+4t+4=100 (t+2)^2=100 t+2=+10 or -10

jhonyy9 (jhonyy9):

(2x+10)(2x+6)=96 multiplie term by term 2x*2x +2x*6 +10*2x +10*6 =96 4x^2 +12x +20x +60 =96 4x^2 +32x +60 =96 add to both sides (-96) 4x^2 +32x -36 =0 divide both sides by 4 x^2 +8x -9 =0 (x+9)(x-1)=0 x+9=0 x=-9 x-1=0 x=1 so because a length not can being negativ so than x=1 will be the right answer good luck bye